# On the Baltimore Light RailLink into the quantum future

### v1

@@ Appendix C D-Wave quantum annealer details
Table 2: Physical properties of the D-Wave Advantage_system6.3 machine [45].
@@
$\text{M}_{\text{AFM}}$ (pH) [^6] Maximum available mutual inductance achievable between pairs of flux qubit bodies. $1.554$
@@
$\Phi^{i}_{\text{CCJJ}}$ ($\Phi_{0}$) [^7] Initial value of the external flux applied to qubit compound Josephson-junction structures at the start of an anneal. $-0.624$
$\Phi^{f}_{\text{CCJJ}}$ ($\Phi_{0}$)  [^8] Final value at the end of an anneal. $-0.723$
Readout time ($\mu$s) [^9] Typical readout times for reading between one qubit and the full QPU. $18.0-173.0$
Approx. programming Time ($\mu$s) [^10] Typical for problems run on this QPU. Actual problem programming times may vary slightly depending on the nature of the problem. $\sim 14200$
@@
For our experiments with quantum annealing, we made use of the D-Wave Advantage_system6.3 device. All the relevant physical parameters for this device are gathered in Table 2.
This device contains 5614 superconducting qubits which are coupled according to the Pegasus graph structure [34]. Since the native coupling graph of the annealer is quite far from being fully-dense, most problems must be modified to fit on the device through the use of an embedding procedure [46]. For example, one of the scheduling problems with #$\#J=2$ trains and maximal secondary delay $d_{\rm max}=6$ defined in Sec. III is mapped to a QUBO with 42 variables. After translating to an Ising model and embedding into the Pegasus graph, this problem required the use of approximately $80$ to $90$ qubits on the D-Wave machine (the embedding procedure is non-deterministic and small variations are expected). The final embedding for this problem is illustrated in Fig. 13.
Figure 13: Illustration of the embedding of a two train scheduling problem with maximal secondary delay $d_{\max}=6$, requiring $42$ variables into the native graph of the D-Wave annealer, for both disturbed (left) and non-disturbed (right) scenarios. After embedding, roughly 80 qubits were required. Dots represent active qubits and edges active coupling between qubits.
@@ Appendix D IonQ quantum computer details
Table 3: Characterization of the two gate-based quantum processors used in this work: IonQ’s Aria-1 trapped-ion device [47] and IBM’s superconducting Brisbane device [48].
@@
$T_{1}$ $100{\rm s}$ [^11] Average value $231.12\mu{\rm s}$ [^12] Median value
$T_{2}$ $1{\rm s}$\@footnotemark $145.97\mu{\rm s}$\@footnotemark
@@
Single-qubit gate error 0.040%\@footnotemark 0.025%\@footnotemark
@@
Two-qubit gate error 4.250%\@footnotemark 0.742%\@footnotemark
@@
SPAM accuracy 99.520%\@footnotemark
Readout error 1.320%\@footnotemark
As a representative of the gate-based approach to quantum computing, we chose to make use of the IonQ Aria-1 device. This is a gate-based quantum computer with 25 trapped Ytterbium ions acting as qubits, supporting all-to-all connectivity (meaning, two-qubit gates are directly supported between any pair of qubits). The various parameters which characterize the performance of the Aria-1 device are shown in Table 3. Note that the average gate errors are subject to significant variation over time, though they generally appear to be within a factor of 2×2$\times$2 × in either direction of the numbers reported in the table.
@@
Figure 14: A quantum circuit diagram showing the single-layer QAOA ansatz for the simplest 6-variable QUBO, requiring 9 two-qubit gates (ZZ gates couple the outermost qubits the boxes cover) and a circuit depth of 17. This circuit is expressed in terms of an abstract gate set, and must be transpiled into the native gate set for any particular quantum processor for execution. The initial layer of Hadamard gates prepares the ground state of the initial Hamiltonian, then the layer of RZ and ZZ gates corresponds to the evolution under the cost Hamiltonian, and finally the layer of RX gates correspond to the mixer Hamiltonian.
Unlike with the D-Wave machine, since the Aria-1 device has all-to-all connectivity there is no embedding overhead and thus the number of binary variables in the original QUBO is the number of qubits required. For the smallest problem we considered, with one train and $d_{\rm max}=2$, this number is six. The circuit implementing the single-layer QAOA ansatz for this problem is illustrated in Fig. 14. In this circuit, the number of two-qubit gates (depicted as ZZ) required scales with the number of couplings in the Ising model (essentially, the number of off-diagonal non-zero matrix elements in the original $Q$ matrix defining the QUBO). The required circuit depth is not obvious from the figure, as we do not know the actual circuit run on the device after transpilation and so do not know if there are additional barriers or constraints to parallelism introduced at that stage. Nonetheless, we show in Fig. 15 the two-layer QAOA ansatz for the same problem in the abstract gate set. The number of two-qubit gates is exactly doubled since the circuit consists of two copies of the same subcircuit run one after another, however the figure shows that the required circuit depth could be significantly less than double. Since two-qubit gate errors are very likely the dominant error mechanism for the execution of these circuits this has little effect on the overall error, though it does indicate that the runtime need not double.
Figure 15: A quantum circuit diagram showing the two-layer QAOA ansatz for the simplest 6-variable QUBO, again in terms of an abstract gate set. This circuit contains 18 two-qubit gates and has a depth of 25.
@@ Appendix E Results with a superconducting IBM device
Another representative of the gate-based approach to quantum computing is provided by the superconducting quantum computers from IBM. Several of their 127-qubit Eagle r3 devices are available through the IBM Quantum Platform, for example the Brisbane device whose properties are summarized in Table 3.
@@
Unfortunately, due to recent changes to how the IBM Quantum Platform operates and to the requirements which must be satisfied by circuits submitted to the cloud interface, the Qiskit implementation of QAOA we have used in this work can not currently be run on IBMs hardware [49]. Therefore, we can only present results from the noisy simulator which was provided as part of the platform until recently [50]. Using the ibmq_qasm_simulator configured to emulate the noise characteristics of the Brisbane device, we ran each of the three smallest scheduling problems (corresponding to 6, 10, and 14 variable QUBOs) five times, with two choices of penalty values yielding an overlapping or split spectrum. In terms of the abstract gate set, the circuits submitted were essentially the same as the ones used on the Aria-1 device, e.g. as in Fig. 14 for the 6 variable problem. Unlike with the Aria-1 device, however, it is likely that the transpilation process would add additional two-qubit gates over the minimum number set by the $Q$ matrix due to the restricted connectivity of the IBM device. For circuits with as few qubits as the ones we examine this is likely not a serious problem but for larger instances this overhead may dominate.
The distributions of final energies returned from the noisy simulator are plotted in Fig. 16. While it is impossible to draw reliable conclusions from so few data points, comparing these results to the equivalent results from the IonQ Aria-1 simulator seem to indicate that the performance of the IBM device should be expected to be comparable to the IonQ device, assuming that the simulator accurately captures the behavior of the device.
Figure 16: Histograms of the result energies yielded by noisy simulations of the IBM Brisbane superconducting gate-based quantum computer running QAOA with a single-layer ansatz. Each QUBO was minimized five times.

### v2

@@ Appendix B Hybrid algorithm for partially-stochastic railway networks
The proposed hybrid algorithm is presented below in bullet points:
Fixed inputs
  1. 1.
    Problem topology, parameters, constraints;
  2. 2.
    Problem decomposition into stochastic (smaller) component and deterministic (larger) component;
  3. 3.
    Train timetable, optionally including initial disturbances;
  4. 4.
    Particular statistics of the train traffic on the stochastic part;
  5. 5.
    Objective function;
Processing
  1. 1.
    From initial data (timetable, disturbances), define the sub-problem on the stochastic component and encode it in QUBO form;
  2. 2.
    Solve the QUBO on a quantum device;
  3. 3.
    From QUBO solutions select feasible solutions, check the statistics of railway traffic parameters of the output, compare with expected statistics and re-sample or filter if necessary;
  4. 4.
    Use selected QUBO solutions as inputs to the deterministic component encoded as an ILP problem, and run multiple optimizations with multiple inputs;
  5. 5.
    Select a series of feasible solutions of the whole system;
  6. 6.
    Prepare new QUBO for stochastic component given the solution of the stochastic component in point $5$, then repeat points and $2$-$5$ check whether objective improves;
  7. 7.
    Execute point $6$ till stopping condition.
Output
  1. 1.
    Series of solutions with various conflict-free timetables given various possible events in the stochastic component.
@@ Appendix C D-Wave quantum annealer details
Maximum achievable mutual inductance $\text{M}_{\text{AFM}}$ (pH) $1.554$
@@
Initial value of external flux $\Phi^{i}_{\text{CCJJ}}$ ($\Phi_{0}$) $-0.624$
Final value of external flux $\Phi^{f}_{\text{CCJJ}}$ ($\Phi_{0}$) $-0.723$
Typical readout time, one qubit to full QPU ($\mu$s) $18.0-173.0$
Typical programming time ($\mu$s) $\sim 14200$
@@
Table C.1: Physical properties of the D-Wave Advantage_system6.3 machine [58].
For our experiments with quantum annealing, we made use of the D-Wave Advantage_system6.3 device. All the relevant physical parameters for this device are gathered in Table C.1.
This device contains 5614 superconducting qubits which are coupled according to the Pegasus graph structure [39]. Since the native coupling graph of the annealer is quite far from being fully-dense, most problems must be modified to fit on the device through the use of an embedding procedure [59]. For example, one of the scheduling problems with #$\#J=2$ trains and maximal secondary delay $d_{\rm max}=6$ defined in Sec. III is mapped to a QUBO with 42 variables. After translating to an Ising model and embedding into the Pegasus graph, this problem required the use of approximately $80$ to $90$ qubits on the D-Wave machine (the embedding procedure is non-deterministic and small variations are expected). The final embedding for this problem is illustrated in Fig. C.1.
The D-Wave embedding strategy uses chains of physical qubits to represent individual logical qubits. Here, we have used the default setting of dwave.system, the EmbeddingComposite Python library for all calculations on the D-Wave. The actual percentage of the number of solutions with no broken chains is greater than $50\%$ for all calculations. In the worst case (#$\#J=11$ trains, maximal secondary delay $d_{\max}=6$, and annealing time $10\mu$s) we had still $55\%$ of solutions without chain break.
Figure C.1: Illustration of the embedding of a two train scheduling problem with maximal secondary delay $d_{\max}=6$, requiring $42$ variables into the native graph of the D-Wave annealer, for both disturbed (left) and non-disturbed (right) scenarios. After embedding, roughly 80 qubits were required. Dots represent active qubits and edges active coupling between qubits.
@@ Appendix D IonQ quantum computer details
$T_{1}$ $100{\rm s}$ $231.12\mu{\rm s}$
$T_{2}$ $1{\rm s}$ $145.97\mu{\rm s}$
@@
Single-qubit gate error 0.040% 0.025%
@@
Two-qubit gate error 4.250% 0.742%
@@
SPAM accuracy 99.520%
Readout error 1.320%
Table D.1: Characterization of the two gate-based quantum processors used in this work: IonQ’s Aria-1 trapped-ion device [60] and IBM’s superconducting Brisbane device [61]. The reported errors and lifetimes are average values for the Aria-1 device and median values for the Brisbane device.
As a representative of the gate-based approach to quantum computing, we chose to make use of the IonQ Aria-1 device. This is a gate-based quantum computer with 25 trapped Ytterbium ions acting as qubits, supporting all-to-all connectivity (meaning, two-qubit gates are directly supported between any pair of qubits). The various parameters which characterize the performance of the Aria-1 device are shown in Table D.1. Note that the average gate errors are subject to significant variation over time, though they generally appear to be within a factor of 2×2$\times$2 × in either direction of the numbers reported in the table.
@@
Figure D.1: A quantum circuit diagram showing the single-layer QAOA ansatz for the simplest 6-variable QUBO, requiring 9 two-qubit gates (ZZ gates couple the outermost qubits the boxes cover) and a circuit depth of 17. This circuit is expressed in terms of an abstract gate set, and must be transpiled into the native gate set for any particular quantum processor for execution. The initial layer of Hadamard gates prepares the ground state of the initial Hamiltonian, then the layer of RZ and ZZ gates corresponds to the evolution under the cost Hamiltonian, and finally the layer of RX gates correspond to the mixer Hamiltonian.
Unlike with the D-Wave machine, since the Aria-1 device has all-to-all connectivity there is no embedding overhead and thus the number of binary variables in the original QUBO is the number of qubits required. For the smallest problem we considered, with one train and $d_{\rm max}=2$, this number is six. The circuit implementing the single-layer QAOA ansatz for this problem is illustrated in Fig. D.1. In this circuit, the number of two-qubit gates (depicted as ZZ) required scales with the number of couplings in the Ising model (essentially, the number of off-diagonal non-zero matrix elements in the original $Q$ matrix defining the QUBO). The required circuit depth is not obvious from the figure, as we do not know the actual circuit run on the device after transpilation and so do not know if there are additional barriers or constraints to parallelism introduced at that stage. Nonetheless, we show in Fig. D.2 the two-layer QAOA ansatz for the same problem in the abstract gate set. The number of two-qubit gates is exactly doubled since the circuit consists of two copies of the same subcircuit run one after another, however the figure shows that the required circuit depth could be significantly less than double. Since two-qubit gate errors are very likely the dominant error mechanism for the execution of these circuits this has little effect on the overall error, though it does indicate that the runtime need not double.
Figure D.2: A quantum circuit diagram showing the two-layer QAOA ansatz for the simplest 6-variable QUBO, again in terms of an abstract gate set. This circuit contains 18 two-qubit gates and has a depth of 25.
@@ Appendix E Results with a superconducting IBM device
Another representative of the gate-based approach to quantum computing is provided by the superconducting quantum computers from IBM. Several of their 127-qubit Eagle r3 devices are available through the IBM Quantum Platform, for example the Brisbane device whose properties are summarized in Table D.1.
@@
Unfortunately, due to recent changes to how the IBM Quantum Platform operates and to the requirements which must be satisfied by circuits submitted to the cloud interface, the Qiskit implementation of QAOA we have used in this work can not currently be run on IBMs hardware [62]. Therefore, we can only present results from the noisy simulator which was provided as part of the platform until recently [63]. Using the ibmq_qasm_simulator configured to emulate the noise characteristics of the Brisbane device, we ran each of the three smallest scheduling problems (corresponding to 6, 10, and 14 variable QUBOs) five times, with two choices of penalty values yielding an overlapping or split spectrum. In terms of the abstract gate set, the circuits submitted were essentially the same as the ones used on the Aria-1 device, e.g. as in Fig. D.1 for the 6 variable problem. Unlike with the Aria-1 device, however, it is likely that the transpilation process would add additional two-qubit gates over the minimum number set by the $Q$ matrix due to the restricted connectivity of the IBM device. For circuits with as few qubits as the ones we examine this is likely not a serious problem but for larger instances this overhead may dominate.
The distributions of final energies returned from the noisy simulator are plotted in Fig. E.1. While it is impossible to draw reliable conclusions from so few data points, comparing these results to the equivalent results from the IonQ Aria-1 simulator seem to indicate that the performance of the IBM device should be expected to be comparable to the IonQ device, assuming that the simulator accurately captures the behavior of the device.
Figure E.1: Histograms of the result energies yielded by noisy simulations of the IBM Brisbane superconducting gate-based quantum computer running QAOA with a single-layer ansatz. Each QUBO was minimized five times.